# From Uncertainty to Trust: Enhancing Reliability in Vision-Language Models with Uncertainty-Guided Dropout Decoding

### v1

@@ Abstract
Large vision-language models (LVLMs) demonstrate remarkable capabilities in multimodal tasks but are prone to misinterpreting visual inputs, often resulting in hallucinations and unreliable outputs. To address these challenges, we propose Dropout Decoding, a novel inference-time approach that quantifies the uncertainty of visual tokens and selectively masks uncertain tokens to improve decoding. Our method measures the uncertainty of each visual token by projecting it onto the text space and decomposing it into aleatoric and epistemic components. Specifically, we focus on epistemic uncertainty, which captures perception-related errors more effectively. Inspired by dropout regularization, we introduce uncertainty-guided token dropout, which applies the dropout principle to input visual tokens instead of model parameters, and during inference rather than training. By aggregating predictions from an ensemble of masked decoding contexts, Dropout Decoding robustly mitigates errors arising from visual token misinterpretations. Evaluations on benchmarks including CHAIR, THRONE, and MMBench demonstrate that Dropout Decoding significantly reduces object hallucinations (OH) and enhances both reliability and quality of LVLM outputs across diverse visual contexts. Code is released at https://github.com/kigb/DropoutDecoding.
@@ 1 Introduction
Recent advancements in large vision-language models (LVLMs) have demonstrated impressive capabilities [11, 54, 66, 15, 59, 62, 17], in tasks such as image captioning [58], visual question answering (VQA) [1, 22, 61, 29], multimodal reasoning [37, 63, 40, 6, 39] and so on. However, LVLMs still face challenges in accurately perceiving and interpreting visual inputs, leading to inaccurate outputs and hallucinations [35]. These issues often stem from LVLMs misrepresenting key image elements or overlooking critical details, compromising the reliability of their outputs in tasks demanding precise visual understanding [16, 57, 4, 5].
@@
In practice, LVLMs typically process visual inputs token by token, which we refer to as visual tokens. [^1] We specifically refer to the tokens that are already in the input prompt to the text decoder. Concrete definition is in §3.1. This can fall short in effectively focusing on the most informative parts of the visual context. While attention mechanisms are designed to prioritize relevant information, they are not always perfect [52, 46], especially when the inputs are complex or ambiguous for the model, or in other words, of high uncertainty. Existing methods to address these challenges in the training stage often involve fine-tuning on specific tasks [56, 34, 33, 49], or using additional supervision signals especially at lower level to guide the model [55, 7]. However, these approaches are resource-intensive and not easily extensible to new tasks. Alternative inference-time strategies, such as attention-based or logits-based mechanisms on decoding correction [51, 45, 4, 21, 60], attempt to identify important regions in the input without additional training, but they typically rely on heuristic design choices and largely increase inference cost. Therefore, enhancing the trustworthiness of LVLMs and reducing hallucinations require more principled methods that can more effectively emphasize the most informative parts of the visual input.
@@
To address this challenge, we propose a novel approach that quantifies uncertainty in visual token contexts and removes uncertain tokens, both directly at inference time to improve the reliability of LVLM outputs. Inspired by traditional dropout  [47] techniques—typically applied to model parameters but difficult to implement directly in pretrained LVLMs [13, 24]—we introduce token dropout, which applies the dropout principle to input context tokens instead of model parameters. Furthermore, it is applied to regularize the inference process instead of training, by introducing randomness in decoding contexts to reduce overfitting to noisy visual tokens.
@@
Our method measures the uncertainty of each visual token by projecting it into the text token space through the text decoder directly, and decomposing this uncertainty into two components: aleatoric (data-related) and epistemic (model-related) [50, 20, 43]. By focusing on epistemic uncertainty, which reflects the model’s lack of knowledge, we identify visual tokens with high uncertainty and selectively target them for suppression. At inference time, we adjust the visual inputs by selectively suppressing tokens with high epistemic uncertainty. Specifically, we create an ensemble of predictions by generating multiple subsets of visual inputs, each with different combinations of high-uncertainty tokens dropped out. These subsets are processed independently, and their corresponding outputs are aggregated using majority voting to produce the final prediction.
This approach, which we term Dropout Decoding, enhances the reliability and accuracy of LVLM outputs without modifying the underlying model parameters or requiring additional training. By leveraging uncertainty quantification and token dropout, Dropout Decoding robustly mitigates errors arising from uncertain visual token interpretations. Experiments are conducted on LVLM decoding benchmarks including CHAIR [41], THRONE [23], and MMBench [36], demonstrating the effectiveness of our approach in both reducing object hallucinations (OH) and improving reliability of model outputs consistently across diverse visual contexts. With Dropout Decoding, we make the following contributions:
  • We introduce a novel approach that quantifies and decomposes uncertainty on tokens in the visual inputs at inference time without additional supervision, by projecting visual input tokens onto text token interpretations.
  • We propose a decoding strategy that uses epistemic uncertainty measurements to guide the selective dropout of high-uncertainty visual tokens in the context, analogous to performing dropout on the model but applied to the input tokens and during inference.
  • Comprehensive experiments are conducted on various benchmarks, showing significant reductions in OH and improved fidelity in pre-trained LVLMs without additional fine-tuning.
@@ 2 Related Work
Reliable generation.
Reliable generation in LLMs is often challenged by hallucinations, where the model generates irrelevant or factually incorrect information [18, 64, 48]. These hallucinations stem from issues in data, training, and inference stages [53], with attention mechanisms exacerbating the problem as sequence lengths grow [8]. To mitigate these, methods like factual-nucleus sampling have been proposed to balance output diversity and accuracy [26]. Besides, while Arias et al. [2] leverage quantified uncertainty to guide the decoding process for LLM, our method differs significantly. We quantify uncertainty at the level of visual input context rather than of model ensemble which is heavy.
OH in LVLMs.
OH is a common issue in LVLMs, where models generate descriptions containing objects, attributes, or relationships not present in the actual image. The CHAIR metric [41] is widely used to evaluate OH, measuring the hallucination rate on the MSCOCO dataset [31]. Another benchmark, POPE [30], treats object hallucination as a binary classification task. More recently, THRONE [23] takes a more holistic approach, using open-ended, object-based image descriptions for evaluation. In our work, we use CHAIR and THRONE to assess OH.
OH reduction. Recent methods addressing OH in LVLMs include internal signal guidance, contrastive decoding, and selective information focusing, all of which are inference-time strategies. OPERA [19] uses internal signals like attention patterns to refine outputs and improve alignment with visual content. Contrastive decoding methods, like VCD [27], enhance coherence by comparing image-specific outputs. Selective information focusing approaches, such as HALC [4], prioritize key image regions, while CDG [10] uses CLIP embeddings to align generation with visual input. In contrast, Dropout Decoding works with any LVLM by 1) selecting visual information from visual tokens during generation, unlike HALC which selects regions initially; 2) using uncertainty to guide visual information selection, requiring no external models, unlike HALC and CDG; 3) introducing a token-level majority voting strategy.
@@ 3.1 Vision-Language Model Decoding
Widely adopted LVLM architectures [28, 32, 33] typically include a vision encoder, a vision-text interface module, and a Transformer-based LLM decoder. As we mostly focus on the decoder side inference optimization, we assume the LLM decoder is with parameter $\theta$.
The visual input, such as an image, is segmented into patches and processed by the vision encoder, [^2] We assume a general Transformer architecture for the vision encoder as well. Our approach could also apply to other types of vision encoders. followed by the vision-text interface module, to produce a sequence of visual tokens $x^{v}=(x^{v}_{1},x^{v}_{2},\dots,x^{v}_{N})$. Each token $x^{v}_{i}$ is a contextualized embedding of an image patch, serving as the direct input to the text decoder. The text input such as a query or instruction is $x^{t}=(x^{t}_{1},x^{t}_{2},\dots,x^{t}_{M})$. The input to the text decoder is denoted as $x=[x^{v},x^{t}]$, which is the concatenation of visual and text tokens. At this point, the visual and text tokens are aligned and serve as a sequential input to the LLM decoder.
During autoregressive decoding, the decoder generates output text tokens $y=(y_{1},y_{2},\ldots)$ as continuation from prompt $x$, following the conditional probability distribution
$h_{j}=$ $\;f_{\theta}(x^{v},x^{t},y_{<j})$ (1)
$p_{\theta}(y_{j}\mid x^{v},x^{t},y_{<j})=$ $\;\mathrm{softmax}(W_{\mathcal{V}}h_{j})$
@@ 3.2 Uncertainty Quantification
Our approach quantifies the information uncertainty of visual tokens used for decoding by adapting the concept of epistemic uncertainty for measurement, as detailed in §5, and drawing inspiration from classical uncertainty decomposition [20, 44, 43]. To provide the necessary background, we first introduce the concept of uncertainty decomposition.
@@
$=\underbrace{\mathbb{E}_{p(w\mid\mathcal{D})}\left[\mathbf{\mathbb{H}}(p(y\mid x,w))\right]}_{\text{Aleatoric Uncertainty}}$
$+\underbrace{\mathbb{E}_{p(w\mid\mathcal{D})}\left[D_{\mathrm{KL}}(p(y\mid x,w)\parallel p(y\mid x,\mathcal{D}))\right]}_{\text{Epistemic Uncertainty}}$
The epistemic uncertainty, expressed as the KL divergence between candidate models’ predictive distributions and the BMA, has proven effective in various applications [38, 3, 13, 65]. Our approach, adopts a similar formulation for uncertainty quantification, calculating the KL divergence between candidate prediction distributions on individual visual tokens and an aggregated average distribution.
@@ 4 Textual Interpretation of Visual Tokens
As discussed in §1, identifying the visual tokens that carry significant information and quantifying their uncertainty is critical for improving the reliability of LVLMs. To address this, we propose a supervision-free, scalable approach that maps visual tokens to the text token space, effectively translating visual content into an interpretable text-based representation. This mapping acts as a heuristic for understanding visual tokens, leveraging the LVLM’s inherent ability to align visual and textual contexts.
@@
Building on this intuition, we adopt a heuristic approach to interpret visual tokens by projecting them onto the text vocabulary at the top Transformer layers. In particular, for each visual token $x^{v}_{i}$ at position $i$, [^4] Note that $i$ indexes are only used over visual tokens $x^{v}$, not text tokens $x^{t}$ or generations $y$. we obtain its textual projected distribution over the vocabulary $\mathcal{V}$ from the last layer of the LLM decoder in the LVLM as:
@@
where $h^{v}_{i}$ is the LLM decoder top-layer hidden representation aligned at the $i$-th visual token positions, $x^{v}_{\leq i}$ denotes the visual tokens up until index $i$ [^5] For the models we use, the visual tokens $x^{v}$ are all placed before the text tokens $x^{t}$ in the concatenated sequence $x$, so $x^{v}_{\leq i}$ are purely visual tokens. But our approach also applies to other cases.
@@
Here, $q^{\text{proj}}_{i}$, which we refer to as visual-textual distribution, represents the projection of the visual input onto the text space. It encapsulates the model’s interpretation of the $i$-th visual token. This projection offers a text-based summarization, akin to an unordered caption or a “bag-of-words” representation of the visual content. As we will demonstrate in §6, this heuristic method serves as an effective proxy for uncertainty estimation.
@@
Figure 1: An illustrative example where visual tokens are projected into the text space. We show 5 image patches and their corresponding top-5 projected text tokens. Words in bold indicate high information content. Red text highlights projections that are evidently misaligned with the image content. This demonstrates that the visual tokens sometimes cannot perfectly capture the image’s information, which motivates the token dropout (§5.2). We also show the uncertainty values quantified on top (dotted lines are average among all visual tokens for reference). High epistemic uncertainty correlates well with high information in visual tokens, whereas aleatoric and total uncertainty do not.
An illustrative example with projection uncertainty. To illustrate the effectiveness of this projection method and motivate our approach, consider the example shown in Fig. 1. The image is processed into patches, and for five selected patches, we compute their corresponding distribution over the text space. Then we obtain the top-$5$ predicted text tokens for each.
Some patches produce specific and informative text tokens, often corresponding to meaningful visual content like “Berlin,” “computer,” or “map.” These tokens are relatively closer to the long tail in vocabulary, indicating that the corresponding visual tokens capture unique and informative visual context. In contrast, patches resulting in common words carry less specific information, because high-frequency words (e.g. “a,” “the,” or “on”) contribute less to the uniqueness of the visual content. This observation suggests that the projected text tokens can serve as a proxy for the information content of the visual tokens.
Building on these insights, we introduce uncertainty measures derived from textual projective distributions $q^{\text{proj}}_{i}$, which we use in our method to quantify the uncertainty associated with each visual token, as shown at the top of Fig. 1. Specifically, inspired by §3.2, we decompose the total uncertainty into two components: aleatoric (data-related), calculated directly from $q^{\text{proj}}_{i}$, and epistemic (model-related), obtained by comparing $q^{\text{proj}}_{i}$ to an average distribution. Detailed definitions are in §5.1.
As shown in Fig. 1, the epistemic uncertainty accurately reflects the information content of the visual tokens: visual tokens with high epistemic uncertainty correspond to patches with significant information (e.g., “Berlin”), while those with low epistemic uncertainty correspond to less informative patches (e.g., “the”). In contrast, the aleatoric and the total uncertainty do not correlate well. This finding motivates our focus on epistemic uncertainty as a reliable indicator of the significance of visual information.

### v2

# Enhancing Vision-Language Model Reliability with Uncertainty-Guided Dropout Decoding

@@ Abstract
Large vision-language models (LVLMs) excel at multimodal tasks but are prone to misinterpreting visual inputs, often resulting in hallucinations and unreliable outputs. We present Dropout Decoding, a novel inference-time approach that quantifies the uncertainty of visual tokens and selectively masks uncertain tokens to improve decoding. Our method measures the uncertainty of each visual token by projecting it onto the text space and decomposing it into aleatoric and epistemic components. Specifically, we focus on epistemic uncertainty, which captures perception-related errors more effectively. Inspired by dropout regularization, we introduce uncertainty-guided token dropout, which applies the dropout principle to input visual tokens instead of model parameters, and during inference rather than training. By aggregating predictions from an ensemble of masked decoding contexts, we can robustly mitigate errors arising from visual token misinterpretations. Evaluations on benchmarks including CHAIR, THRONE, and MMBench demonstrate that Dropout Decoding significantly reduces object hallucinations (OH) and enhances both reliability and quality of LVLM outputs across diverse visual contexts. Code is released at https://github.com/kigb/DropoutDecoding.
@@ 1 Introduction
Recent advancements in large vision-language models (LVLMs) have demonstrated impressive capabilities (du2022survey; li20254d; zhao2024multimodal) in tasks such as image captioning, visual question answering (VQA), and multimodal reasoning (antol2015vqa; ma2023crepe; gupta2023cliptrans; li2024socialgpt). However, LVLMs still face challenges in accurately perceiving and interpreting visual inputs, leading to inaccurate outputs and hallucinations (liu2024survey). These issues often stem from LVLMs misrepresenting key image elements or overlooking critical details (gunjal2023detecting; zhai2023halle). In practice, LVLMs typically process visual inputs token by token (li2025-textOrPixels), which we refer to as visual tokens. [^1] We specifically refer to the tokens that are already in the input prompt to the text decoder. Concrete definition is in §3.1. This can fall short in effectively focusing on the most informative parts of the visual context. While attention mechanisms are designed to prioritize relevant information, they are not always perfect (selvaraju2020grad; attentionrag; kangsee), especially when the inputs are complex or ambiguous for the model, or in other words, of high uncertainty. Existing methods to address these challenges in the training stage often involve fine-tuning on specific tasks (yuksekgonul2023visionlanguagemodelsbehavelike; liu2024llavanext; liu2024llava1_5; wang2024mpo_internvl; song2025head), or using additional supervision signals especially at lower level to guide the model (you2023ferret_ground_anything; cheng2024spatialrgpt). However, these approaches are resource-intensive and not easily extensible to new tasks. Alternative inference-time strategies rely on attention or logits-based mechanisms but typically use heuristic designs and increase inference cost (xu2015show; selvaraju2017grad; chen2024halcobjecthallucinationreduction; zhang2024sled_self_logits_decoding). Therefore, enhancing the trustworthiness of LVLMs (jiang2025robustifying) and reducing hallucinations (chen2024halcobjecthallucinationreduction) require more principled methods that can more effectively emphasize the most informative parts of the visual input.
@@
To address this challenge, we propose a novel approach that quantifies uncertainty in visual token contexts and removes uncertain tokens, both directly at inference time to improve the reliability of LVLM outputs. Inspired by traditional dropout (srivastava2014dropout) techniques—typically applied to model parameters but difficult to implement directly in pretrained LVLMs (gal2016dropout; kendall2017uncertainties)—we introduce token dropout, which applies the dropout principle to input context tokens instead of model parameters. Furthermore, it is applied to regularize the inference process instead of training, by introducing randomness in decoding contexts to reduce overfitting to noisy visual tokens.
@@
Our method measures the uncertainty of each visual token by projecting it into the text token space through the text decoder directly, and decomposing this uncertainty into two components: aleatoric (data-related) and epistemic (model-related) (wilson2020bayesian; hullermeier2021aleatoric; schweighofer2023introducingimprovedinformationtheoreticmeasure). By focusing on epistemic uncertainty, which reflects the model’s lack of knowledge, we identify visual tokens with high uncertainty and selectively target them for suppression. At inference time, we adjust the visual inputs by selectively suppressing tokens with high epistemic uncertainty. Specifically, we create an ensemble of predictions by generating multiple subsets of visual inputs, each with different combinations of high-uncertainty tokens dropped out. These subsets are processed independently, and their corresponding outputs are aggregated using majority voting to produce the final prediction.
Our method, termed Dropout Decoding, enhances the reliability and accuracy of LVLM outputs without modifying the underlying model parameters or requiring additional training. Experiments on LVLM decoding benchmarks including CHAIR (rohrbach2019objecthallucinationimagecaptioning), THRONE (kaul2024throneobjectbasedhallucinationbenchmark), and MMBench (liu2024mmbenchmultimodalmodelallaround) demonstrate the effectiveness of our approach. In summary, we make the following contributions. First, we introduce a novel approach that quantifies and decomposes uncertainty on tokens in the visual inputs at inference time without additional supervision, by projecting visual input tokens onto text token interpretations. Second, we propose a decoding strategy that uses epistemic uncertainty measurements to guide the selective dropout of high-uncertainty visual tokens in the context, analogous to performing dropout on the model but applied to the input tokens and during inference. And finally, comprehensive experiments are conducted on various benchmarks, showing significant reductions in OH and improved fidelity in pre-trained LVLMs without additional fine-tuning.
@@ 2 Related Work
Reliable Generation. Hallucinations in LLMs—where models generate irrelevant or incorrect information (huang2023surveyhallucinationlargelanguage; zhao2024enhanced; wang2024preference)—arise from data (song2025head), training, and inference issues (xu2024hallucinationinevitableinnatelimitation), with attention mechanisms exacerbating them (chiang-cholak-2022-overcoming). To address this, factual-nucleus sampling (lee2023factualityenhancedlanguagemodels) balances diversity and accuracy. While (arias2024adaptive_contrastive_uncertainty_guided_decoding) guide decoding with quantified uncertainty, our approach quantifies uncertainty at the visual input level, not requiring model ensembles.
OH in LVLMs. Object hallucination (OH) is common in LVLMs, where models generate incorrect object descriptions. CHAIR (rohrbach2019objecthallucinationimagecaptioning) and POPE (li2023evaluatingobjecthallucinationlarge) evaluate OH, while THRONE (kaul2024throneobjectbasedhallucinationbenchmark) offers a more holistic approach. We use CHAIR and THRONE to assess OH in our work.
OH Reduction. Methods addressing OH in LVLMs include internal signal guidance (e.g., OPERA (huang2024operaalleviatinghallucinationmultimodal)), contrastive decoding (e.g., VCD leng2023mitigatingobjecthallucinationslarge), and selective information focusing (e.g., HALC (chen2024halcobjecthallucinationreduction)). AGLA (an2024aglamitigatingobjecthallucinations) mitigates hallucinations by enhancing visual grounding through global and local attention, while Memory-Space Visual Retracing (zou2024looktwiceanswermemoryspace) refines multimodal alignment via iterative visual reference retrieval. In contrast, Dropout Decoding 1) selects visual tokens during generation, 2) uses uncertainty for token selection without external models, and 3) employs a token-level majority voting strategy.
@@ 3.1 Vision-Language Model Decoding
Widely adopted LVLM architectures li2022blip; liu2023llava; liu2024llava1_5 typically include a vision encoder, a vision-text interface module, and a Transformer-based LLM decoder. As we mostly focus on the decoder side inference, we assume the LLM decoder parameterized by $\theta$.
The visual input, such as an image, is segmented into patches and processed by the vision encoder, [^2] We assume a general Transformer architecture for the vision encoder as well. Our approach could also apply to other types of vision encoders. followed by the vision-text interface module, to produce a sequence of visual tokens $x^{v}=(x^{v}_{1},x^{v}_{2},\dots,x^{v}_{N})$. Each token $x^{v}_{i}$ is a contextualized embedding of an image patch, serving as the direct input to the text decoder. The text input such as a query or instruction is $x^{t}=(x^{t}_{1},x^{t}_{2},\dots,x^{t}_{M})$. The input to the text decoder is denoted as $x=[x^{v},x^{t}]$, which is the concatenation of visual and text tokens. At this point, the visual and text tokens are aligned and serve as a sequential input to the LLM decoder. During autoregressive decoding, the decoder generates output text tokens $y=(y_{1},y_{2},\ldots)$ as continuation from prompt $x$, following the conditional probability distribution
$h_{j}=\;f_{\theta}(x^{v},x^{t},y_{<j}),\quad p_{\theta}(y_{j}\mid x^{v},x^{t},y_{<j})=\;\mathrm{softmax}(W_{\mathcal{V}}h_{j})$ (1)
@@ 3.2 Uncertainty Quantification
Our approach quantifies the information uncertainty of visual tokens used for decoding by adapting the concept of epistemic uncertainty for measurement, as detailed in §5, and drawing inspiration from classical uncertainty decomposition (hullermeier2021aleatoric; schweighofer2023quantification; schweighofer2023introducingimprovedinformationtheoreticmeasure). To provide the necessary background, we first introduce the concept of uncertainty decomposition.
@@
$=\underbrace{\mathbb{E}_{p(w\mid\mathcal{D})}\left[\mathbf{\mathbb{H}}(p(y\mid x,w))\right]}_{\text{Aleatoric Uncertainty}}+\underbrace{\mathbb{E}_{p(w\mid\mathcal{D})}\left[D_{\mathrm{KL}}(p(y\mid x,w)\parallel p(y\mid x,\mathcal{D}))\right]}_{\text{Epistemic Uncertainty}}$
The epistemic uncertainty, expressed as the KL divergence between candidate models’ predictive distributions and the BMA, has proven effective in various applications (osband2016deepexplorationbootstrappeddqn; burda2018explorationrandomnetworkdistillation; gal2016dropout; zhao2024direct). Our approach, adopts a similar formulation for uncertainty quantification, calculating the KL divergence between candidate prediction distributions on individual visual tokens and an aggregated average distribution.
@@ 4 Textual Interpretation of Visual Tokens
As discussed in §1, identifying the visual tokens that carry significant information and quantifying their uncertainty is critical for improving the reliability of LVLMs. We propose a supervision-free approach that maps visual tokens to text token space for improving LVLM reliability by identifying significant visual tokens and quantifying their uncertainty. This mapping leverages the LVLM’s inherent ability to align visual and textual contexts.
@@
Building on this intuition, we adopt a heuristic approach to interpret visual tokens by projecting them onto the text vocabulary at the top Transformer layers. In particular, for each visual token $x^{v}_{i}$ at position ii, [^4] Note that $i$ indexes are only used over visual tokens $x^{v}$, not text tokens $x^{t}$ or generations $y$. we obtain its textual projected distribution over the vocabulary $\mathcal{V}$ from the last layer of the LLM decoder in the LVLM as:
@@
where $h^{v}_{i}$ is the LLM decoder top-layer hidden representation aligned at the $i$-th visual token positions, $x^{v}_{\leq i}$ denotes the visual tokens up until index ii. [^5] For the models we use, the visual tokens $x^{v}$ are all placed before the text tokens $x^{t}$ in the concatenated sequence $x$, so $x^{v}_{\leq i}$ are purely visual tokens. But our approach also applies to other cases. This approach is also generally referred to as logit lens (belrose2023elicitinglatentpredictionstransformers) in mechanistic interpretability for LLMs.
@@
Here, $q^{\text{proj}}_{i}$, which we refer to as visual-textual distribution, represents the projection of the visual input onto the text space. It encapsulates the model’s interpretation of the $i$-th visual token. This projection offers a text-based summarization, akin to an unordered caption or a “bag-of-words” (yuksekgonul2023visionlanguagemodelsbehavelike) representation of the visual content. As we will demonstrate in §6, this heuristic method serves as an effective proxy for uncertainty estimation.
@@
Figure 1: An illustrative example where visual tokens are projected into the text space, bold words indicate highly informative projections, and red words mark misalignments. Dotted lines show average uncertainties; high epistemic uncertainty correlates with informative patches.
An illustrative example with projection uncertainty. Figure 1 demonstrates our projection method by processing an image into patches and projecting five selected patches into the text space, retrieving their top-5 text tokens. Informative patches yield specific tokens like “Berlin,” “computer,” or “map,” which are less frequent in the vocabulary and capture unique visual contexts. In contrast, patches producing common words (e.g., “a,” “the,” “on”) convey less specific information. This suggests that projected text tokens effectively proxy the information content of visual tokens.
Leveraging this, we introduce uncertainty measures from the textual projection distributions $q^{\text{proj}}_{i}$ to quantify each visual token’s uncertainty, as depicted in the figure. Following classical uncertainty quantification (§3.2), we decompose total uncertainty into aleatoric (data-related) derived directly from $q^{\text{proj}}_{i}$, and epistemic (model-related) by comparing $q^{\text{proj}}_{i}$ to an average distribution (§5.1). As illustrated, epistemic uncertainty aligns well with the information content of visual tokens: high epistemic uncertainty corresponds to informative patches (e.g., “Berlin”), and vice versa (e.g., “the”). In contrast, aleatoric and total uncertainty do not show this correlation. This finding motivates our focus on epistemic uncertainty as a reliable indicator of the significance of visual information.